At STP, 50 mL of isopropanol (C3H7OH, density: 0.786 g/mL) is combusted in the presence of 2 L O2. Which starting material remains after the reaction and what is the remaining volume?
1
Expert's answer
2012-05-29T07:39:04-0400
C3H7OH + 5O2=3CO2+4H2O

m=50 ml*0.786 g/ml=39.3 g
n=m/Mr=39.3/60=0.655
mol

n(O2)=2/22.4=0.0893 mol
oxygen react completely
only
0.0893/5=0.0179 isopropanol react and it remains after reaction
=0.655-0.0179=0.63714 mol

V of CO2 =(0.0893/5 *3)=0.0536*22.4=1.20 L